You have a box that measures 2 meters x 2 meters x 6 meters. What is the longest piece of rebar (inside diagonal) that could fit in this box? What assumption is necessary for this calculation?


1
Expert's answer
2021-08-29T17:37:52-0400

length (l)=2 meters

breadth(b)=2 meters

height(h)=6 meters

the longest piece of rebar(x)

"x=\sqrt{l^2+b^2+h^2}\\x=\sqrt{^2+2^2+6^2}\\x=\sqrt{4+4+36}\\x=\sqrt{44}\\x=2\sqrt{11}"

so the longest piece of rebar ="2\sqrt{11}" meters

for this calculation l, b ,h and x assumption are necessary.